OpenStudy (anonymous):

Which function rule is represented by the table? x 0 1 2 3 y 3 5 7 9 (1 point) y = x + 3 y = x + 6 y = 2x + 3 y = 4x + 1

OpenStudy (cwrw238):

try plugging in values of x into each equation and see if they fit eg x = 0 into y = x + 3 gives y = 0 + 3 = 3 - that one looks good but you need to check more values to be sure

OpenStudy (anonymous):